# Finite Difference Solution Ansatz approach in Least-Squares Monte Carlo

### v7

@@ Abstract
This article presents a simple but effective and efficient approach to improve the accuracy and stability of Least-Squares Monte Carlo for American-style option pricing as well as expected exposure calculation in valuation adjustments. The key idea is to construct the ansatz of conditional expected continuation payoff using the finite difference solution from one dimension, to be used in linear regression. This approach bridges between solving backward partial differential equations and Monte Carlo simulation, aiming at achieving the best of both worlds. Independent of model settings, the ansatz is proved to serve as a control variate to reduce the least-squares errors. We illustrate the technique with realistic examples including Bermudan options, worst of issuer callable notes and expected positive exposure on European options under valuation adjustments. The method can be considered as a generic numerical scheme across various asset classes, in particular, as an accurate method for pricing and risk-managing American-style derivatives under arbitrary dimensions.
@@ 1 Introduction
  • Our method is suitable under generic model settings, not limited to stochastic volatility only.
@@
The rest of this paper is organized as follows. Sec.2 is the description of the framework from a theoretical standpoint, where we first introduce background knowledge in Sec.2.1 in the context of American-style exercise derivatives pricing and CVA. After setting up the general problem in Sec.2.1.1, we outline prevailing numerical methods in practice, including the least-squares Monte Carlo approach in Sec.2.1.2, and finite difference PDE method in Sec.2.1.4. With such foundation, we will proceed to introduce our novel approach in Sec.2.2 from a model-independent assumption, by highlighting its enhancement over the classical approach in Sec.2.2.2 and subsequently performing error analysis in Sec.2.2.3. Then underlying model settings and implementation details will be given in Sec.2.2.4 and Sec.2.2.5, respectively. Numerical results for Bermudan options and worst of issuer callable notes are presented in Sec.3, which examines the accuracy of various available methods. The CVA calculation on European option are also presented. Finally, concluding remarks are given in Sec.4.
@@ 2.1.1 Problem Formulation
As a general framework relevant for derivative valuation as well as regulatory CVA computations, we assume that throughout a given terminal expiry $T$ with the riskless interest rate $r$, there exists a complete probability space ($\Omega,\operatorname{\mathcal{F}},(\operatorname{\mathcal{F}}_{t})_{0\leq t\leq T},\operatorname*{\mathbb{Q}}$) satisfying the usual conditions with respect to which all processes are defined under risk-neutral measure $\operatorname*{\mathbb{Q}}$. Also, let $L^{2}(\Omega,\operatorname{\mathcal{F}},(\operatorname{\mathcal{F}}_{t})_{0\leq t\leq T},\operatorname*{\mathbb{Q}})$, hereinafter simplified as $L^{2}$, denote the space of square integrable functions. As problem formulation, from $t=0$ and final expiry $T$ of an derivative contract, we define a set of discretized time points as $\pi(t):=\{t_{k}\geq t|t_{k}=k\Delta T,1\leq k\leq M=\frac{T}{\Delta T}\}$ with a time interval $\Delta T$. $\pi(t)$ represents either the exercise times for American-style option, or discrete exposure monitoring times in CVA calculation. Within the complete probability space $\Omega$, there exists a discrete filtration $\{\operatorname{\mathcal{F}}(t_{k})|1\leq k\leq M\}$ on $\pi(0)$. For notational clarity, we use $XYZ(\omega,t)$ to denote a stochastic variable under $\operatorname{\mathcal{F}}(t)$ with $\omega\in\Omega$. In this spirit, the state variables are denoted by $X(\omega,t_{k})$, which is $\operatorname{\mathcal{F}}(t_{k})$-measurable.
@@ 2.1.2 Least-Squares Monte Carlo
As can be seen in Eqn.2.1, once the $\operatorname{\mathcal{F}}_{t_{k}}$-measurable conditional expectations can be efficiently crystallized into a functional form depending on $X(\omega,t_{k})$, the PV or CVA can be determined in a path-by-path manner independently and thus estimated by its corresponding Monte Carlo estimators. The core idea of regression based Least-Squares approach (LSM) is to approximate the parameterized form of $\operatorname{\mathbb{E}}[Y(\omega,t)|\operatorname{\mathcal{F}}_{t_{k}}]$ (omitting the superscript of $Y$) as a linear combination of a collection of $R$ basis functions $\vec{\phi}^{P}_{R}(x)=\{\phi^{P}_{r}(x)|r=1,\ldots,R\}$ in $\operatorname*{\mathbb{R}}^{R}$ space, where the coefficients are obtained via regression using the cross-sectional information computed in the simulation. For notational convenience, let’s define the least-squares estimator of the conditional expectation, using the $L^{2}$ projection operator $P_{k}$ at $t_{k}$ under $\vec{\phi}^{P}_{R}(x)$ , as
@@
$\begin{split}\vec{\beta}^{P}_{k}&=\arg_{\vec{b}\in\mathbb{R}^{R}}\min\mathcal{L}^{P}(\vec{b};Y(\omega,t)),\\ \mathcal{L}^{P}(\vec{b};Y(\omega,t))&=\operatorname{\mathbb{E}}\left[(Y(\omega,t)-\vec{b}\cdot\vec{\phi}^{P}_{R}(X(\omega,t_{k})))^{2}\right|\operatorname{\mathcal{F}}_{0}].\end{split}$ (2.3)
It means that $\vec{\beta}^{P}_{k}=(\beta^{P}_{k,1},\ldots,\beta^{P}_{k,R})^{T}$ are to be obtained via regression as described in Appendix A.3. In Eqn.2.3, $\mathcal{L}^{P}(\vec{b};Y(\omega,t))$ is the objective function to be minimized, whose lower bound is given by, in the limit of $\vec{b}\cdot\vec{\phi}^{P}_{R}(X(\omega,t_{k}))$ being exact:
$\inf_{\vec{b}\in\mathbb{R}^{R}}\mathcal{L}^{P}(\vec{b};Y(\omega,t))=\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(Y(\omega,t)|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}],$ (2.4)
@@
In this article, we study Bermudan options on constant-weighted basket $S_{B}(t)=\sum_{i=1}^{d}w_{i}S_{i}(t)$ with $w_{i}=\frac{1}{d}$ (see Appendix A.1), and worst of issuer callable notes (WIC) on worst-of basket $S_{W}(t)=\min_{i\in\{1,\ldots,d\}}S_{i}(t)$ (see Appendix A.2), respectively. Here the underlying basket is $d$ dimensional and $\vec{S}(t)=(S_{1}(t),\ldots,S_{d}(t))^{T}$ are asset spot prices.
Under discrete-time formulation, a preferable way to solve the Eqn.2.5 is to find out the optimal stopping time strategy, denoted by $\tau(t_{k}):=\tau(\omega,t_{k})$ (removing the tilde compared to $\tilde{\tau}(t_{k})$ for optimality and omitting the dependence on $\omega$ for notational convenience). For further notational clarity, we also introduce $V(\omega,\tau(t_{k+1}))$ to represent the option path value under optimal stopping time $\tau(t_{k})$, differentiated from $\tilde{V}(\omega,\tilde{\tau}(t_{k+1}))$ by removing the tilde similarly to emphasize that it is calculated under an optimal stopping policy.
@@
Given the solution of Eqn.2.7, we arrive at the desired result
@@
which will be useful in the theoretical analysis of pricing error by translating a global optimization problem in Eqn.2.5 into a dynamic programming problem.
To proceed under LSM following Eqn.2.2, we introduce the regressed continuation value
@@
In practice, we run two Monte Carlo simulations, where in the first regression stage $F_{k}^{P}(X(\omega,t_{k}))$ is determined, and then PV is evaluated in the subsequent pricing stage. It is worth mentioning that in the regression stage, when solving $\vec{\beta}^{P}_{k}$ iteratively using Eqn.2.7 and Eqn.2.2, 2.3, with $e^{-r\Delta T}V(\omega,\tau(t_{k+1}))$ substituted for $Y$ ($Y^{PV}$) as the regression dependent variable, the least-squares estimator only enters into the calculation inside the exercise indicator $I(\omega,t_{k})$, whereas the continued value function on the right hand side of Eqn.2.7 is still calculated as pathwise cash flows from the simulation. Stentoft, (2014) showed that the approach by Tsitsiklis and Van Roy, (2001), which further replaces the value function outside the exercise indicator outright with its estimator, results in a much higher absolute bias.
@@
Now we turn to the example of CVA computation, as the expected loss resulting from the potential future default of the counterparty. For the ease of illustration without loss of generality, we assume that the product doesn’t have early termination features, e.g., an European Option on $S_{B}(T)$. By choosing $Y^{CVA}(\omega,T):=\tilde{V}(\omega,\tilde{\tau}(t_{k})=T)=e^{-r(T-t_{k})}Z(S_{B}(T))$ [^1] One can consider the European option as a limiting case of Bermudan option by forcing all $\tilde{\tau}(t_{k})=T$. whose conditional expectation $\tilde{F}(\omega,t_{k}):=\operatorname{\mathbb{E}}[Y^{CVA}(\omega,T)|{\operatorname{\mathcal{F}}}_{t_{k}}]$ is the mark-to-market value of the option as of $t_{k}$, the CVA on this financial transaction can be derived, in the form of Eqn.2.1, as:
@@
where $R_{ctpy}$ is the recovery rate and $PD(t)$ denotes the default probability of the counterparty at $t$, which depends on the counterparty hazard rate denoted by $h_{cpty}(\omega,t)$.
@@
To further take into account the Wrong Way Risk (WWR), we consider a hazard rate with a general functional dependence on the option mark-to-market value $\tilde{F}(\omega,t)$(Hull and White, (2012)):
$h_{cpty}(\omega,t)\equiv h_{cpty}(\tilde{F}(\omega,t),t)=\ln[1+e^{a(t)+b\tilde{F}(\omega,t)}],$
where $a(t)$ and $b$ are WWR model parameters.
@@
respectively. Here the superscript $(i)$ indicates that the variable is under $i^{th}$ of the total $N_{P}$ pricing paths, similar to PV problem in Eqn.2.9.
It is worth mentioning that in Eqn.2.14 and 2.15, the value function are completely approximated with its regressed estimator. Admittedly, this approach is more akin to the procedure advocated by Tsitsiklis and Van Roy, (2001), which could result in biases that are difficult to control without using pathwise cash flows calculation in pricing stage; however, it is still the standard market practice for the indispensable sake of efficient trade aggregation at the netting set level.  [^2] Note that we are aware of alternative hybrid approach to calculate CVA/EPE, by making use of both the simulated cash flows and regressed function (demonstrated by Joshi and Kwon, (2016) without considering WWR). The drawback of this proposal is that one has to generate all the cash flows at given individual product level, without being decoupled from the path generation at the upper portfolio level. It would result in an aggregation of contractual timelines, rather than simple algorithmic operation upon regressed functions at individual product level when it comes to portfolio aggregation.

### v8

@@ Abstract
This article presents a simple but effective and efficient approach to improve the accuracy and stability of Least-Squares Monte Carlo. The key idea is to construct the ansatz of conditional expected continuation payoff using the finite difference solution from one dimension, to be used in linear regression. This approach bridges between solving backward partial differential equations and Monte Carlo simulation, aiming at achieving the best of both worlds. In a general setting encompassing both local and stochastic volatility models, the ansatz is proven to act as a control variate, reducing the mean squared error, thereby leading to a reduction of the final pricing error. We illustrate the technique with realistic examples including Bermudan options, worst of issuer callable notes and expected positive exposure on European options under valuation adjustments.
@@ 1 Introduction
  • Our method is suitable under a wider range of model settings, not limited to stochastic volatility only.
@@
The rest of this paper is organized as follows. Sec.2 is the description of the framework from a theoretical standpoint, where the underlying model settings are presented in Sec.2.1. Then we introduce background knowledge in Sec.2.2 in the context of American-style exercise derivatives pricing and CVA. After setting up the general problem in Sec.2.2.1, we outline prevailing numerical methods in practice, including the least-squares Monte Carlo approach (Sec.2.2.2 and its limitation in Sec.2.2.3), and finite difference PDE method (Sec.2.2.4). With such foundation, and further motivated by the analysis of a limiting case in Sec.2.3.1, we proceed to introduce our novel approach in Sec.2.3, by highlighting its enhancement over the classical approach in Sec.2.3.2 and subsequently performing error analysis in Sec.2.3.3. Then implementation details will be given in Sec.2.3.4. The numerical results are delivered in Section 3, which compares the new method against existing alternatives regarding accuracy, efficiency, and stability: Sec.3.1, 3.2, and 3.3 examine Bermudan option pricing under the Black-Scholes (up to 4D) and Heston models; Sec.3.4 details CVA calculations for European options under 4D; last but not least, Sec.3.5 presents examples of worst-of issuer callable notes under a multi-variate local volatility model. Finally, concluding remarks are given in Sec.4.
@@ 2 Theoretical Framework
2.1 Model Settings
For model settings on the underlying diffusion process, we consider multi-variate local volatility and single-asset stochastic volatility. In either case, an equivalent one-dimensional process (often a local volatility model) exists that matches the marginal distributions of the original process, as demonstrated by Gyöngy, (1986).
Multi-variate Local Volatility
for $i=1\ldots d$, where $q_{i}$ and $\sigma_{i}(S_{i}(t),t)$ are the constant continuous dividend yield and local volatility of the $i$-th asset’s motion, respectively; and $\{B_{i}(t)|i=1\ldots d\}$ is a correlated $d$-dimensional Brownian motion, with $\operatorname{\mathbb{E}}[dB_{i}(t)dB_{j}(t)]=\rho_{ij}dt:=[\delta_{ij}+\rho(1-\delta_{ij})]dt$ using Kronecker delta notation $\delta_{ij}$ and $\rho$ denoting single correlation between assets. Since we only consider continuous (rather than discrete) dividends, we assume $\vec{S}(0)=(1,\ldots,1)^{T}$ for simplicity without loss of generality. In the limit of $\sigma_{i}(S_{i}(t),t)$ being constants, this setting degenerates to the celebrated Black-Scholes model.
Stochastic Volatility: Heston Model
For stochastic volatility setting, let’s consider below stochastic differential equation under the Heston model
with $\operatorname{\mathbb{E}}[dB^{S}(t)dB^{\nu}(t)]=\rho^{S,\nu}dt$, where $q$ is the constant continuous dividend yield, $\theta$ is the long variance, $\kappa$ is the mean reverting rate, $\xi$ is the volatility of volatility and $\rho^{S,\nu}$ is the correlation between spot and variance. For the same reason as previous subsection, we assume $S(0)=1$ for simplicity. And the initial variance is given by $\nu(0)$. There is a well known result on the expectation value of $\nu(t)$ over time: $\operatorname{\mathbb{E}}[\nu(t)]=\nu(0)e^{-\kappa t}+\theta(1-e^{-\kappa t})$.
@@ 2.2.1 Problem Formulation
As a general framework relevant for derivative valuation as well as regulatory CVA computations, we assume that throughout a given terminal expiry $T$ with the riskless interest rate $r$ defined previously, there exists a complete probability space ($\Omega,\operatorname{\mathcal{F}},(\operatorname{\mathcal{F}}_{t})_{0\leq t\leq T},\operatorname*{\mathbb{Q}}$), accommodating the model settings in Sec.2.1, under risk-neutral measure $\operatorname*{\mathbb{Q}}$. Also, let $L^{2}(\Omega,\operatorname{\mathcal{F}},(\operatorname{\mathcal{F}}_{t})_{0\leq t\leq T},\operatorname*{\mathbb{Q}})$, hereinafter simplified as $L^{2}$, denote the space of square integrable functions. As problem formulation, from $t=0$ and final expiry $T$ of an derivative contract, we define a set of discretized time points as $\pi(t):=\{t_{k}\geq t|t_{k}=k\Delta T,1\leq k\leq M=\frac{T}{\Delta T}\}$ with a time interval $\Delta T$. $\pi(t)$ represents either the exercise times for American-style option, or discrete exposure monitoring times in CVA calculation. Within the complete probability space $\Omega$, there exists a discrete filtration $\{\operatorname{\mathcal{F}}(t_{k})|1\leq k\leq M\}$ on $\pi(0)$. For notational clarity, we use $XYZ(\omega,t)$ to denote a stochastic variable under $\operatorname{\mathcal{F}}(t)$ with $\omega\in\Omega$. In this spirit, the state variables are denoted by $X(\omega,t_{k})$, which is $\operatorname{\mathcal{F}}(t_{k})$-measurable.
@@ 2.2.2 Least-Squares Monte Carlo
As can be seen in Eqn.2.3, once the $\operatorname{\mathcal{F}}_{t_{k}}$-measurable conditional expectations can be efficiently crystallized into a functional form depending on $X(\omega,t_{k})$, the PV or CVA can be determined in a path-by-path manner independently and thus estimated by its corresponding Monte Carlo estimators. The core idea of regression based Least-Squares approach (LSM) is to approximate the parameterized form of $\operatorname{\mathbb{E}}[Y(\omega,t)|\operatorname{\mathcal{F}}_{t_{k}}]$ (omitting the superscript of $Y$) as a linear combination of a collection of $R$ basis functions $\vec{\phi}^{P}_{R}(x)=\{\phi^{P}_{r}(x)|r=1,\ldots,R\}^{T}$ in $\operatorname*{\mathbb{R}}^{R}$ space, where the coefficients are obtained via regression using the cross-sectional information computed in the simulation. For notational convenience, let’s define the least-squares estimator of the conditional expectation, using the $L^{2}$ projection operator $P_{k}$ at $t_{k}$ under $\vec{\phi}^{P}_{R}(x)$, as
@@
$\begin{split}\vec{\beta}^{P}_{k}=\arg_{\vec{b}\in\mathbb{R}^{R}}\min\mathcal{L}(Y(\omega,t);\vec{b};\vec{\phi}^{P}_{R}(X(\omega,t_{k}))),\\ \mathcal{L}(Y(\omega,t);\vec{b};\vec{\phi}^{P}_{R}(X(\omega,t_{k}))):=\operatorname{\mathbb{E}}\left[(Y(\omega,t)-\vec{b}\cdot\vec{\phi}^{P}_{R}(X(\omega,t_{k})))^{2}\right|\operatorname{\mathcal{F}}_{0}].\end{split}$ (2.5)
$P_{k}$ is a projector onto a linear span of $L^{2}$, defined by a given basis functions $\vec{\phi}^{P}_{R}(X(\omega,t_{k}))$ as
$\text{span}(\vec{\phi}^{P}_{R}(X(\omega,t_{k}))):=\{\vec{b}\cdot\vec{\phi}^{P}_{R}(X(\omega,t_{k}))|\vec{b}\in{\operatorname*{\mathbb{R}}}^{R}\}.$ (2.6)
In Eqn.2.5, $\mathcal{L}(Y(\omega,t);\vec{b};\vec{\phi}^{P}_{R}(X(\omega,t_{k})))$ is the objective function to be minimized, which is the residual sum of squares (RSS). Its lower bound is given by, in the limit of $\vec{b}\cdot\vec{\phi}^{P}_{R}(X(\omega,t_{k}))\to\operatorname{\mathbb{E}}[Y(\omega,t)|X(\omega,t_{k})]$,
$\inf_{\vec{b}\in\mathbb{R}^{R}}\mathcal{L}(Y(\omega,t);\vec{b};\vec{\phi}^{P}_{R}(X(\omega,t_{k})))=\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(Y(\omega,t)|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}],$ (2.7)
@@
The formal solution of Eqn.2.5 as a function of $Y(\omega,t)$ is given by
$\vec{\beta}^{P}_{k}(Y(\omega,t)):=(\operatorname{\mathbb{E}}[\vec{\phi}^{P}_{R}(X(\omega,t_{k}))\vec{\phi}^{P}_{R}(X(\omega,t_{k}))^{T}|\operatorname{\mathcal{F}}_{0}])^{-1}\operatorname{\mathbb{E}}[\vec{\phi}^{P}_{R}(X(\omega,t_{k}))Y(\omega,t)|\operatorname{\mathcal{F}}_{0}].$ (2.8)
In practice, the Monte Carlo estimator of $\vec{\beta}^{P}_{k}$ can be obtained via regression as described in Eqn.4.1 of Appendix A.3.
@@
In this article, we study Bermudan options on constant-weighted basket $S_{B}(t)=\sum_{i=1}^{d}w_{i}S_{i}(t)$ with $w_{i}=\frac{1}{d}$ (see Appendix A.1), and worst of issuer callable notes (WIC) on worst-of basket $S_{W}(t)=\min_{i\in\{1,\ldots,d\}}S_{i}(t)$ (see Appendix A.2), respectively.
Under discrete-time formulation, a preferable way to solve the Eqn.2.9 is to find out the optimal stopping time strategy, denoted by $\tau(t_{k}):=\tau(\omega,t_{k})$ (removing the tilde compared to $\tilde{\tau}(t_{k})$ for optimality and omitting the dependence on $\omega$ for notational convenience). For further notational clarity, we also introduce $V(\omega,\tau(t_{k+1}))$ to represent the option path value under optimal stopping time $\tau(t_{k})$, differentiated from $\tilde{V}(\omega,\tilde{\tau}(t_{k+1}))$ by removing the tilde similarly to emphasize that it is calculated under an optimal stopping policy.
@@
Given the solution of Eqn.2.11, we arrive at the desired result
@@
which will be useful in the theoretical analysis of pricing error by translating a global optimization problem in Eqn.2.9 into a dynamic programming problem.
To proceed under LSM following Eqn.2.4, we introduce the regressed continuation value
@@
In practice, we run two Monte Carlo simulations, where in the first regression stage $F_{k}^{P}(X(\omega,t_{k}))$ is determined, and then PV is evaluated in the subsequent pricing stage. It is worth mentioning that in the regression stage, when solving $\vec{\beta}^{P}_{k}$ iteratively using Eqn.2.11 and Eqn.2.4, 2.5, with $e^{-r\Delta T}V(\omega,\tau(t_{k+1}))$ substituted for $Y$ ($Y^{PV}$) as the regression dependent variable, the least-squares estimator only enters into the calculation inside the exercise indicator $I(\omega,t_{k})$, whereas the continued value function on the right-hand side of Eqn.2.11 is still calculated as pathwise cash flows from the simulation. Stentoft, (2014) showed that the approach by Tsitsiklis and Van Roy, (2001), which further replaces the value function outside the exercise indicator outright with its estimator, results in a much higher absolute bias.
@@
Now we turn to the example of CVA computation, as the expected loss resulting from the potential future default of the counterparty. For the ease of illustration without loss of generality, we assume that the product doesn’t have early termination features, e.g., an European Option on $S_{B}(T)$. By choosing $Y^{CVA}(\omega,T):=\tilde{V}(\omega,\tilde{\tau}(t_{k})=T)=e^{-r(T-t_{k})}Z(S_{B}(T))$ [^1] One can consider the European option as a limiting case of Bermudan option by forcing all $\tilde{\tau}(t_{k})=T$. whose conditional expectation $\tilde{F}(\omega,t_{k}):=\operatorname{\mathbb{E}}[Y^{CVA}(\omega,T)|{\operatorname{\mathcal{F}}}_{t_{k}}]$ is the mark-to-market value of the option as of $t_{k}$, the CVA on this financial transaction can be derived, in the form of Eqn.2.3, as:
@@
where $R_{ctpy}$ is the recovery rate and $PD(t)$ denotes the default probability of the counterparty at $t$, which depends on the counterparty hazard rate denoted by $h_{cpty}(\omega,t)\equiv h_{cpty}(\tilde{F}(\omega,t),t)$.
@@
respectively. Here the superscript $(i)$ indicates that the variable is under $i^{th}$ of the total $N_{P}$ pricing paths, similar to PV problem in Eqn.2.13.
It is worth mentioning that in Eqn.2.18 and 2.19, the value function are completely approximated with its regressed estimator. Admittedly, this approach is more akin to the procedure advocated by Tsitsiklis and Van Roy, (2001), which could result in biases that are difficult to control without using pathwise cash flows calculation in pricing stage; however, it is still the standard market practice for the indispensable sake of efficient trade aggregation at the netting set level.  [^2] Note that we are aware of alternative hybrid approach to calculate CVA/EPE, by making use of both the simulated cash flows and regressed function (demonstrated by Joshi and Kwon, (2016) without considering Wrong Way Risk). The drawback of this proposal is that one has to generate all the cash flows at given individual product level, without being decoupled from the path generation at the upper portfolio level. It would result in an aggregation of contractual timelines, rather than simple algorithmic operation upon regressed functions at individual product level when it comes to portfolio aggregation.